OpenStudy (anikay):

If I'm given: Vertices at (-2,5) and (-2,1) Foci (-2,6) and (-2,0) then how do I find the equation of the hyperbole?

OpenStudy (mertsj):

If you know the vertices and foci you can determine the orientation of the hyperbola. So do that first. Does it open right and left or up and down?

OpenStudy (anikay):

up and down

OpenStudy (mertsj):

So the center is on the line x = -2 half way between the vertices. What is the center?

OpenStudy (anikay):

(-2,3) right?

OpenStudy (mertsj):

Yes. And since it opens up and down, we know that y is the positive term. Do you agree?

OpenStudy (anikay):

yes. absolutely

OpenStudy (mertsj):

Hang on a sec. Just had an interruption.

OpenStudy (anikay):

of course :)

OpenStudy (mertsj):

Sorry.

OpenStudy (mertsj):

So now we can write the numerators of our two fractions because we know the center. Can you do that?

OpenStudy (anikay):

one sec, this is like a refresher for me and a new subject for the person I'm with, so I'm going to talk her through this next part, just a second

OpenStudy (mertsj):

ok

OpenStudy (anikay):

The numerators are \[3y^2 and -2x^2\] correct?

OpenStudy (mertsj):

\[\frac{(y-k)^2}{a^2}-\frac{(x-h)^2}{b^2}=1\]

OpenStudy (mertsj):

(h,k) is the center

OpenStudy (anikay):

oh okay, that's the best standard equation I've found all day. I love this site.

OpenStudy (mertsj):

So now, if we can just find "a" and "b" we'll be done.

OpenStudy (mertsj):

And remember a is the distance from the center to the vertex.

OpenStudy (anikay):

b is vertex to focus correct?

OpenStudy (mertsj):

b is the distance from the center to the co-vertex (which we don't know) and c is the distance from the center to the focal point. Fortunately we know that a, b, and c are related by this equation: a^2+b^2=c^2

OpenStudy (mertsj):

So what is a? What is c?

OpenStudy (anikay):

9-4 so b is sqrt(5)

OpenStudy (anikay):

uh a^2 is 9? thought it was 4

OpenStudy (anikay):

oh you meant c

OpenStudy (mertsj):

2^2+b^2=3^2

OpenStudy (anikay):

\[\frac{ (y-3)^2 }{ 4 } -\frac{ (x+2) }{ 5 } = 3\] is what I got

OpenStudy (mertsj):

it's all good except the right side should be =1

OpenStudy (anikay):

okay, thank you :) big help, it's always 1?

OpenStudy (mertsj):

Yes. When the equation is in the "h,k" form

OpenStudy (anikay):

okay thanks

OpenStudy (mertsj):

yw

OpenStudy (anikay):

if it opens left right, A still goes under the first term right?

OpenStudy (mertsj):

Yes. a^2 under positive term.

OpenStudy (anikay):

ty

OpenStudy (mertsj):

yw